A company is analyzing two mutually exclusive projects, S and L, with the following cash flows:

0 1 2 3 4
Project S -$1,000 $877.56 $260 $10 $5
Project L -$1,000 $5 $240 $400 $808.14

The company's WACC is 10.5%. What is the IRR of the better project? (Hint: The better project may or may not be the one with the higher IRR.) Round your answer to two decimal places.
Answer #1

S:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=877.56/1.105+260/1.105^2+10/1.105^3+5/1.105^4

=$1017.87

NPV=Present value of inflows-Present value of outflows

=1017.87-1000

=$17.87(Approx).

L:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=5/1.105+240/1.105^2+400/1.105^3+808.14/1.105^4

=$1039.59

NPV=Present value of inflows-Present value of outflows

=1039.59-1000

=$39.59(Approx).

Hence L is better having higher NPV.

Let irr be x%
At irr,present value of inflows=present value of outflows.

1000=5/1.0x+240/1.0x^2+400/1.0x^3+808.14/1.0x^4

Hence x=irr=11.8%(Approx).